OpenStudy (anonymous):

2m/m-9 + 18/9-m? 2m/m-9 + 18/9-m? @Mathematics

OpenStudy (anonymous):

Do I do it like this: 2m 18 --- + ----- m-9 9-m 9-m = -(-9 +m) = -(m-9)

OpenStudy (anonymous):

2m 18 --- + ----- m-9 -(m-9) = 2m 18 ---- - ----- m-9 m-9

OpenStudy (anonymous):

and I'm stuck D: